# Reality conditions for the KdV equation and quasi-periodic solutions in finite phase spaces

### v2

@@ Abstract.
In this paper reality conditions for finite-gap solutions of the KdV equation are determined. This new result brings the last piece of the puzzle of obtaining quasi-periodic solutions of completely integrable dynamical equations. Effective computation of the solutions is suggested, and illustrated in genera $2$ and $3$.
Real-valued quasi-periodic solutions, which describe nonlinear waves, are proposed in every finite phase space of the hierarchy of hamiltonian systems of the KdV equation. These solutions are expressed in terms of the abelian $\wp_{1,1}$-function, which comes as a result of algebro-geometric integration. The construction of the hierarchy on coadjoint orbits of an $\mathrm{SL}(2,\operatorname{\mathbb{R}})$ loop group is also addressed.
@@ 1. Introduction
The Korteweg—de Vries equation (KdV) arose in the $19$th century in connection with the theory of waves in shallow water [^1] The KdV equation was mentioned in the footnote on page 360 in Boussinesq, J., Essai sur la theorie des eaux courantes, Memoires presentes par divers savants, l’Acad. des Sci. Inst. Nat. France, XXIII (1877), pp. 1–680. However, D. J. Korteweg and G. de Vries, (1895) gave the full explanation in [15]., see [16]. It is known that this equation also describes the propagation of waves with weak dispersion in various nonlinear media, see [15]. The conventional form of KdV is
(1) $w_{t}=6ww_{x}+w_{xxx}.$
@@
This equation is scale-invariant, that is multiplying $t$, $x$, and $w$ by constants can be used to make the coefficients of any of the three terms equal to any given non-zero constants.
@@
First known solutions: the one-soliton solution and the simplest non-linear wave solution — were suggested by Korteweg and de Vries in [16].
@@
In [13] a remarkable procedure of finding soliton solutions of KdV, known as the inverse scattering method, was discovered. Soon, it was shown that the KdV equation admits the Lax representation [19], and possesses a sequence of integrals of motion [21], which tends to be infinite. Further, in [11], it was proven, that the sequence of integrals of motion of KdV is infinite, and so KdV was called a completely integrable hierarchy of hamiltonian systems. In [22] the notion of a finite-gap solution of KdV arose; such a solution lives within a hamiltonian system of a fixed dimension. Also the higher KdV equations were introduced in [22].
Wide variety of integrable equations on orbits of coadjoint representations of finite Lie groups was introduced in [20, 12]. A further development of the orbit approach was given in [1, 2, 17]
@@
The paper is organized as follows. In Preliminaries we recall all notions related to uniformization of hyperelliptic curves: the standard not normalized differentials of the first kind and associated to them differentials of the second kind, Abel’s map, theta and sigma functions, the Jacobi inversion problem, and also briefly explain the construction of hamiltonian systems on coadjoint orbits of a loop group. In section 3 hamiltonian systems which constitute the KdV hierarchy are developed. Section 4 is devoted to separation of variables. In section 5 the algebro-geometric integration is explained in application to the KdV hierarchy. Finally, section 6 presents new results on finding real-valued quasi-periodic solutions of the KdV equation in any finite phase space. Section 7 illustrates an effective computation of the quasi-periodic solutions in genera $1$, $2$, and $3$.
@@ 2.1. Hyperelliptic curves
(2) $-y^{2}+z^{2g+1}+\sum_{i=1}^{2g}\lambda_{2i+2}z^{2g-i}=0.$
Let $(e_{i},0)$, $i=1$, …, $12g+$ 1, be finite branch points of the curve (2). In what follows, we denote branch points simply by $e_{i}$. Homology basis is defined after H. Baker [3, § 200]. One can imagine a continuous path through all branch points, which starts at infinity and ends at infinity, see the orange line on fig. 1.
@@
Let $\mathrm{d}u=(\mathrm{d}u_{1},\mathrm{d}u_{3},\dots,\mathrm{d}u_{2g-1})^{t}$ be not normalized differentials of the first kind, and $\mathrm{d}r=(\mathrm{d}r_{1},\mathrm{d}r_{3},\dots,\mathrm{d}r_{2g-1})^{t}$ be differentials of the second kind associated with the first kind differentials, see [3, § 138] for more detail. Actually,
(3a) $\mathrm{d}u_{2n-1}=\frac{z^{g-n}\mathrm{d}z}{\partial_{y}f(z,y)},\quad n=1,\dots,g,$
(3b) $\mathrm{d}r_{2n-1}=\frac{\mathrm{d}z}{\partial_{y}f(z,y)}\bigg{(}z^{g+n-1}+\sum_{j=1}^{2n-2}(2n-1-j)\lambda_{2j}z^{g+n-1-j}\bigg{)},$
where $\partial_{w}f(z,y)=-2y$. The indices of $\mathrm{d}u_{2n-1}$ display the orders of zeros, and the indices of $\mathrm{d}r_{2n-1}$ display the orders of poles.
@@
$\omega_{k}=\oint_{\mathfrak{a}_{k}}\mathrm{d}u,\qquad\qquad\omega^{\prime}_{k}=\oint_{\mathfrak{b}_{k}}\mathrm{d}u.$
@@ 2.2. Abel’s map
The Abel map of a positive divisor $D=\sum_{i=1}^{n}(z_{i},y_{i})$ is defined by
$\mathcal{A}(D)=\sum_{i=1}^{n}\int_{\infty}^{(z_{i},y_{i})}\mathrm{d}u.$
@@
2.3. Jacobi inversion problem
A solution of the Jacobi inversion problem on a hyperelliptic curve is proposed in [3, Art. 216], see also [7, Theorem 2.2]. Let $u=\mathcal{A}(D)$ be the Abel image of a non-special positive divisor $D\in\mathcal{V}^{g}$. Then $D$ is uniquely defined by the system of equations
(4a) $\mathcal{R}_{2g}(z,y;u)\equiv z^{g}-\sum_{i=1}^{g}z^{g-i}\wp_{1,2i-1}(u)=0,$
(4b) $\mathcal{R}_{2g+1}(z,y;u)\equiv 2y+\sum_{i=1}^{g}z^{g-i}\wp_{1,1,2i-1}(u)=0.$
@@ 2.4. Theta function
$\theta(v;\tau)=\sum_{n\in\operatorname{\mathbb{Z}}^{g}}\exp\big{(}\imath\pi n^{t}\tau n+2\imath\pi n^{t}v\big{)}.$
@@
$\theta[\varepsilon](v;\tau)=\exp\big{(}\imath\pi(\varepsilon^{\prime}{}^{t}/2)\tau(\varepsilon^{\prime}/2)+2\imath\pi(v+\varepsilon/2)^{t}\varepsilon^{\prime}/2\big{)}\times\\ \times\theta(v+\varepsilon/2+\tau\varepsilon^{\prime}/2;\tau)$
@@ 2.6. Characteristics and partitions
Let $\mathcal{S}=\{0,1,2,\dots,2g+1\}$ be the set of indices of all branch points of a hyperellipitic curve of genus $g$, and 00 stands for the branch point at infinity. According to [3, § 202] all half-period characteristics are represented by partitions of $\mathcal{S}$ of the form $\mathcal{I}_{\mathfrak{m}}\cup\mathcal{J}_{\mathfrak{m}}$ with $\mathcal{I}_{\mathfrak{m}}=\{i_{1},\,\dots,\,i_{g+1-2\mathfrak{m}}\}$ and $\mathcal{J}_{\mathfrak{m}}=\{j_{1},\,\dots,\,j_{g+1+2\mathfrak{m}}\}$, where $\mathfrak{m}$ runs from 00 to $[(g+1)/2]$, and []delimited-[][$\cdot$][ ⋅ ] denotes the integer part. Index 00, corresponding to infinity, is usually omitted in sets, it is also omitted in computation of cardinality of a set.
@@
$\mathcal{A}(\mathcal{I})=\sum_{i\in\mathcal{I}}\mathcal{A}(e_{i})=\omega\Big{(}\frac{1}{2}\varepsilon(\mathcal{I})+\frac{1}{2}\tau\varepsilon^{\prime}(\mathcal{I})\Big{)}.$
@@
of $\mathcal{A}(\mathcal{I})+K$, where $K$ denotes the vector of Riemann constants. Characteristic $[K]$ of the vector of Riemann constants equals the sum of all odd characteristics of branch points, there are $g$ such characteristics, see [3, § 200, 202]. In the basis of canonical cycles introduced by fig. 1 we have
@@ 3. Integrable systems on coadjoint orbits of a loop group
$H=\frac{1}{2}\begin{pmatrix}1&0\\ 0&-1\end{pmatrix},\quad X=\begin{pmatrix}0&1\\ 0&0\end{pmatrix},\quad Y=\begin{pmatrix}0&0\\ 1&0\end{pmatrix}.$
@@
$\mathfrak{d}=2z\frac{\mathrm{d}}{\mathrm{d}z}+\operatorname{ad}_{H},$
where $\operatorname{ad}$ denotes the adjoint operator in $\mathfrak{g}$, that is $\forall X\in\mathfrak{g}$ $\operatorname{ad}_{H}X=[H,X]$.
Let $\{X_{2m-1},\,Y_{2m-1},\,H_{2m},\mid m\in\operatorname{\mathbb{Z}}\}$, such that
$X_{2m-1}=z^{m-1}X,\quad Y_{2m-1}=z^{m}Y,\quad H_{2m}=z^{m}H,$
form a basis of $\widetilde{\mathfrak{g}}$. In general, an element of the basis will be denoted by $Z_{a,\ell}$, where $a=1$, $2$, $3$, and $\ell$ indicates the degree of $Z_{a,\ell}$, namely $\mathfrak{d}Z_{a,\ell}=\ell Z_{a,\ell}$. Actually, $Z_{1,\ell}=H_{\ell}$, $Z_{2,\ell}=Y_{\ell}$, $Z_{3,\ell}=X_{\ell}$. Let $\mathfrak{g}_{\ell}$ denote the degree $\ell$ eigenspace of $\mathfrak{d}$. Thus, $\mathfrak{g}_{2m-1}=\operatorname{span}\{X_{2m-1},\,Y_{2m-1}\}$, and $\mathfrak{g}_{2m}=\operatorname{span}\{H_{2m}\}$.
@@
$X_{2m-1}^{\ast}\leftrightarrow X_{-2m-1},\quad Y_{2m-1}^{\ast}\leftrightarrow Y_{-2m-1},\quad H_{2m}^{\ast}\leftrightarrow H_{-2m-2},$
where $\{X_{2m-1}^{\ast},\,Y_{2m-1}^{\ast},\,H_{2m}^{\ast},\mid m\in\operatorname{\mathbb{Z}}\}$, such that
$X_{2m-1}^{\ast}=z^{m}X^{\ast},\quad Y_{2m-1}^{\ast}=z^{m-1}Y^{\ast},\quad H_{2m}^{\ast}=z^{m}H^{\ast},$
@@
$H^{\ast}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix},\quad X^{\ast}=\begin{pmatrix}0&0\\ 1&0\end{pmatrix},\quad Y^{\ast}=\begin{pmatrix}0&1\\ 0&0\end{pmatrix}.$
@@
Note, that $\mathfrak{g}_{-1}^{\ast}=\mathfrak{g}_{-1}$, and $\mathfrak{g}_{\ell}^{\ast}$ is dual to $\mathfrak{g}_{-\ell-2}$, if $\ell>1$.

### v4

# Reality conditions for the KdV equation and exact quasi-periodic solutions in finite phase spaces

@@ Abstract.
In the present paper reality conditions for quasi-periodic solutions of the KdV equation are determined completely. As a result, solutions in the form of non-linear waves can be plotted and investigated.
The full scope of obtaining finite-gap solutions of the KdV equation is presented. It is proven that the multiply periodic $\wp_{1,1}$-function on the Jacobian variety of a hyperelliptic curve of arbitrary genus serves as the finite-gap solution, the genus coincides with the number of gaps. The subspace of the Jacobian variety where $\wp_{1,1}$, as well as other Weierstrass-$\wp$-functions, are bounded and real-valued is found in any genus. This result covers every finite phase space of the KdV hierarchy, and can be extended to other completely integrable equations. A method of effective computation of this type of solutions is suggested, and illustrated in genera $2$ and $3$.
@@ 1. Introduction
The Korteweg—de Vries equation (KdV) arose in the $19$th century in connection with the theory of waves in shallow water [^1] The KdV equation was mentioned in the footnote on page 360 in Boussinesq, J., Essai sur la theorie des eaux courantes, Memoires presentes par divers savants, l’Acad. des Sci. Inst. Nat. France, XXIII (1877), pp. 1–680. However, D. J. Korteweg and G. de Vries, (1895) gave the full explanation in [25]., see [25]. The equation also describes the propagation of waves with weak dispersion in various nonlinear media, see [24]. The conventional form of KdV is
(1) $w_{\mathrm{t}}=6ww_{\mathrm{x}}+w_{\mathrm{x}\mathrm{x}\mathrm{x}}.$
@@
This equation is scale-invariant, that is by scaling $\mathrm{t}$, $\mathrm{x}$, and $w$ one can change the coefficients of the three terms arbitrarily.
@@
The first two solutions: the one-soliton solution, and the simplest non-linear wave solution — were found by Korteweg and de Vries in [25].
@@
In [19] a remarkable procedure of finding solutions of KdV, known as the inverse scattering method, was discovered. Although an elegant form of solutions was suggested, the problem was merely transformed into the Gel’fand—Levitan integral equation. Soon, it was shown that the KdV equation admits the Lax representation [28], and possesses a sequence of integrals of motion [32], which tends to be infinite. In [15] it was proven, that the sequence of integrals of motion is infinite, and so KdV was called a completely integrable hierarchy of hamiltonian systems. The notion of a finite-gap solution arose in [33]; such a solution lives on a finite-dimesional phase space. And within the infinite hierarchy, a solution in any dimension can be constructed. Higher KdV equations were also introduced in [33].
In [21] a method of constructing multi-soliton solutions, known as Hirota’s method, was suggested. Soon after that, these multi-soliton solutions were obtained by the inverse scattering method in the case of no reflection of incoming waves [20].
An alternative method for solving the KdV equation, and a wide variety of other integrable equations, comes from the theory of semi-simple Lie groups. An integrable equation arises within a hierarchy of hamiltonian systems on orbits of coadjoint representation of a loop group, see in [1, 31, 26, 2, 16] how the method was developed. This method is known as the orbit approach. It leads to algebraic integration, and produces solutions in terms of functions which uniformize the spectral curve of a hamiltonian system in question.
A finite-gap solution of the KdV equation was suggested in [23]:
(2) $w(\mathrm{x},\mathrm{t})=2\partial_{\mathrm{x}}^{2}\log\theta(\bm{U}\mathrm{x}+\bm{W}\mathrm{t}+\bm{D})+2c,$
in connection with the spectral problem for the Schrödinger operator. The constant $c$ and constant vectors $\bm{U}$, $\bm{W}$, $\bm{D}$ were not specified therein. In [14] a procedure of reconstructing $\bm{U}$ and $\bm{W}$ was suggested, while the constant $c$ was omitted. A solution in the form (2) is defined more accurately in [6, p. 65–66].
Another form of an exact finite-gap solution, see [9], is given by a multiply periodic $\wp_{1,1}$-function, defined on the Jacobian variety of a hyperelliptic curve of any genus, the genus coincides with the number of gaps. The latter solution sheds light on constant quantities in (2), see Remark 7 for more details.
A solution in the form of $\wp_{1,1}$-function is complex-valued and satisfies (1) with arbitrary complex $\mathrm{t}$ and $\mathrm{x}$. Until now, real conditions remained an open problem for this type of solutions. In the present paper the problem of real conditions for $\wp_{1,1}$-solution is solved completely. Simultaneously, the locus in the Jacobian variety where Weierstrass-$\wp$-functions are bounded and real-valued is found in the case of a curve with all real branch points.
In the recent decades, efficient computation and graphical representation of the KdV solutions have been addressed. In [18] solutions of the KdV and KP equations on hyperelliptic curves of genera $2$, $4$, $6$ are computed in Matlab. Solutions of the form (2) are used, and periods are calculated by expanding the integrands as a series of Chebyshev polynomials, and then integrating the polynomials in an appropriate way. This method of spectral approximation was introduced in [17], as an alternative to the known numerical tools of studying theta-functional solutions. The most popular among the latter is the package algcurves in Maple [13], which calculates characteristic quantities of Riemann surfaces, such as homology basis, not normalized period matrices, and the Riemann period matrix. The most recent development in this direction is a new package of computing theta functions and derivatives in Julia [3]. Therein a review of the existing computational tools is presented.
Solutions of the KdV and mKdV equations in terms of $\wp_{1,1}$-function are computed and graphically represented in [30, 29]. Function $\wp_{1,1}$ is expressed in terms of a divisor on a hyperelliptic curve, which gives a solution of the Jacobi inversion problem. Solving the equation is reduced to computing the Abel image of this divisor by means of Euler’s numerical quadrature. A path in the Jacobian variety is constructed numerically to satisfy the reality condition for the solution. This method avoids computing period matrices.
In the present paper, an analytical method of computing Weierstrass-$\wp$-functions is used, see [8] for more details. The method is based on analytical construction of the Riemann surface of a curve. It produces period matrices of the first and second kinds, required for computation of Weierstrass-$\wp$-functions, and allows to compute the Abel image of any point of the curve. The method was developed to supply Weierstrass-$\wp$-functions with appropriate and convenient computational tools. Wolfram Mathematica, designed for symbolic computation, is used for calculation and graphical representation.
The paper presents the full scope of obtaining finite-gap solutions of the KdV equation. We start with constructing the hierarchy of hamiltonian systems of the KdV equation, then explain algebraic integration in detail, and prove that the $\wp_{1,1}$-function serves as a finite-gap solution in any genus. From the hierarchy we find the accurate relation between the arguments of $\wp_{1,1}$ and variables $\mathrm{t}$ and $\mathrm{x}$. Then we find the domain in the Jacobian variety where $\wp_{1,1}$ is bounded and real-valued. This result is proven in arbitrary genus. Finally, quasi-periodic solutions of the KdV equation in genera $2$ and $3$ are illustrated by plots.
@@
The paper is organized as follows. In Preliminaries we recall all notions related to uniformization of hyperelliptic curves: the standard not normalized differentials of the first kind and associated to them differentials of the second kind, Abel’s map, the theta and sigma functions, the Jacobi inversion problem, and also briefly explain the construction of hamiltonian systems on coadjoint orbits of a loop group. In section 3 hamiltonian systems which form the KdV hierarchy are developed. Section 4 is devoted to separation of variables. In section 5 algebro-geometric integration is explained in application to the KdV hierarchy. Finally, section 6 presents new results on finding bounded and real-valued quasi-periodic solutions of the KdV equation in any finite phase space. In section 7 a method of effective computation of Weierstrass-$\wp$-functions is presented, and quasi-periodic solutions are illustrated in genera $1$, $2$, and $3$.
@@ 2.1. Hyperelliptic curves
(3) $f(x,y)\equiv-y^{2}+x^{2g+1}+\sum_{i=1}^{2g}\lambda_{2i+2}x^{2g-i}=0.$
Let $(e_{i},0)$, $i=1$, …, $12g+$ 1, be finite branch points of the curve (3). In what follows, we denote branch points simply by $e_{i}$. Homology basis is defined after H. Baker [4, § 200]. One can imagine a continuous path through all branch points, which starts at infinity and ends at infinity, see the orange line on fig. 1.
@@
Let $\mathrm{d}u=(\mathrm{d}u_{1},\mathrm{d}u_{3},\dots,\mathrm{d}u_{2g-1})^{t}$ be not normalized differentials of the first kind, and $\mathrm{d}r=(\mathrm{d}r_{1},\mathrm{d}r_{3},\dots,\mathrm{d}r_{2g-1})^{t}$ be differentials of the second kind associated with the first kind differentials, see [4, § 138] for more detail. Actually,
(4a) $\mathrm{d}u_{2n-1}=\frac{x^{g-n}\mathrm{d}z}{\partial_{y}f(x,y)},\quad n=1,\dots,g,$
(4b) $\mathrm{d}r_{2n-1}=\frac{\mathrm{d}x}{\partial_{y}f(x,y)}\bigg{(}x^{g+n-1}+\sum_{j=1}^{2n-2}(2n-1-j)\lambda_{2j}x^{g+n-1-j}\bigg{)},$
where $\partial_{y}f(x,y)=-2y$. The indices of $\mathrm{d}u_{2n-1}$ display the orders of zeros, and the indices of $\mathrm{d}r_{2n-1}$ display the orders of poles.
@@
(5) $\omega_{k}=\oint_{\mathfrak{a}_{k}}\mathrm{d}u,\qquad\qquad\omega^{\prime}_{k}=\oint_{\mathfrak{b}_{k}}\mathrm{d}u.$
@@ 2.2. Abel’s map
The Abel map of a positive divisor $D=\sum_{i=1}^{n}(x_{i},y_{i})$ is defined by
$\mathcal{A}(D)=\sum_{i=1}^{n}\int_{\infty}^{(x_{i},y_{i})}\mathrm{d}u.$
@@ 2.3. Theta function
(6) $\theta(v;\tau)=\sum_{n\in\operatorname{\mathbb{Z}}^{g}}\exp\big{(}\imath\pi n^{t}\tau n+2\imath\pi n^{t}v\big{)}.$
@@
(7) $\theta[\varepsilon](v;\tau)=\exp\big{(}\imath\pi(\varepsilon^{\prime}{}^{t}/2)\tau(\varepsilon^{\prime}/2)+2\imath\pi(v+\varepsilon/2)^{t}\varepsilon^{\prime}/2\big{)}\theta(v+\varepsilon/2+\tau\varepsilon^{\prime}/2;\tau)$
@@ 2.4. Sigma function
2.5. Jacobi inversion problem
A solution of the Jacobi inversion problem on a hyperelliptic curve is proposed in [4, Art. 216], see also [9, Theorem 2.2]. Let $u=\mathcal{A}(D)$ be the Abel image of a non-special positive divisor $D\in\mathcal{V}^{g}$. Then $D$ is uniquely defined by the system of equations
(9a) $\mathcal{R}_{2g}(x,y;u)\equiv x^{g}-\sum_{i=1}^{g}x^{g-i}\wp_{1,2i-1}(u)=0,$
(9b) $\mathcal{R}_{2g+1}(x,y;u)\equiv 2y+\sum_{i=1}^{g}x^{g-i}\wp_{1,1,2i-1}(u)=0.$
Here and below $\mathcal{R}_{n}$ denotes an entire rational function of weight $n$ on the curve.
@@ 2.6. Characteristics and partitions
Let $\mathcal{S}=\{0,1,2,\dots,2g+1\}$ be the set of indices of all branch points of a hyperellipitic curve of genus $g$, and 00 stands for the branch point at infinity. According to [4, § 202] all half-period characteristics are represented by partitions of $\mathcal{S}$ of the form $\mathcal{I}_{\mathfrak{m}}\cup\mathcal{J}_{\mathfrak{m}}$ with $\mathcal{I}_{\mathfrak{m}}=\{i_{1},\,\dots,\,i_{g+1-2\mathfrak{m}}\}$ and $\mathcal{J}_{\mathfrak{m}}=\{j_{1},\,\dots,\,j_{g+1+2\mathfrak{m}}\}$, where $\mathfrak{m}$ runs from 00 to $[(g+1)/2]$, and []delimited-[][$\cdot$][ ⋅ ] denotes the integer part. Index 00, corresponding to infinity, is usually omitted in sets, it is also omitted in computation of cardinality of a set.
@@
$\mathcal{A}(\mathcal{I})=\sum_{i\in\mathcal{I}}\mathcal{A}(e_{i})=\omega\Big{(}\tfrac{1}{2}\varepsilon(\mathcal{I})+\tfrac{1}{2}\tau\varepsilon^{\prime}(\mathcal{I})\Big{)}.$
@@
of $\mathcal{A}(\mathcal{I})+K$, where $K$ denotes the vector of Riemann constants. Characteristic $[K]$ of the vector of Riemann constants equals the sum of all odd characteristics of branch points, there are $g$ such characteristics, see [4, § 200, 202]. In the basis of canonical cycles introduced by fig. 1 we have
@@ 3. Integrable systems on coadjoint orbits of a loop group
$\textsf{H}=\frac{1}{2}\begin{pmatrix}1&0\\ 0&-1\end{pmatrix},\quad\textsf{X}=\begin{pmatrix}0&1\\ 0&0\end{pmatrix},\quad\textsf{Y}=\begin{pmatrix}0&0\\ 1&0\end{pmatrix}.$
@@
$\mathfrak{d}=2z\frac{\mathrm{d}}{\mathrm{d}z}+\operatorname{ad}_{\textsf{H}},$
where $\operatorname{ad}$ denotes the adjoint operator in $\mathfrak{g}$, that is $\forall\,\textsf{X}\in\mathfrak{g}$ $\operatorname{ad}_{\textsf{H}}\textsf{X}=[\textsf{H},\textsf{X}]$.
Let $\{\textsf{X}_{2m-1},\,\textsf{Y}_{2m-1},\,\textsf{H}_{2m},\mid m\in\operatorname{\mathbb{Z}}\}$, such that
$\textsf{X}_{2m-1}=z^{m-1}\textsf{X},\quad\textsf{Y}_{2m-1}=z^{m}\textsf{Y},\quad\textsf{H}_{2m}=z^{m}\textsf{H},$
form a basis of $\widetilde{\mathfrak{g}}$. In general, an element of the basis will be denoted by $\textsf{Z}_{a,\ell}$, where $a=1$, $2$, $3$, and $\ell$ indicates the degree of $\textsf{Z}_{a,\ell}$, namely $\mathfrak{d}\textsf{Z}_{a,\ell}=\ell\textsf{Z}_{a,\ell}$. Actually, $\textsf{Z}_{1,\ell}=\textsf{H}_{\ell}$, $\textsf{Z}_{2,\ell}=\textsf{Y}_{\ell}$, $\textsf{Z}_{3,\ell}=\textsf{X}_{\ell}$. Let $\mathfrak{g}_{\ell}$ denote the degree $\ell$ eigenspace of $\mathfrak{d}$. Thus, $\mathfrak{g}_{2m-1}=\operatorname{span}\{\textsf{X}_{2m-1},\,\textsf{Y}_{2m-1}\}$, and $\mathfrak{g}_{2m}=\operatorname{span}\{\textsf{H}_{2m}\}$.
@@
$\textsf{X}_{2m-1}^{\ast}\leftrightarrow\textsf{X}_{-2m-1},\quad\textsf{Y}_{2m-1}^{\ast}\leftrightarrow\textsf{Y}_{-2m-1},\quad\textsf{H}_{2m}^{\ast}\leftrightarrow\textsf{H}_{-2m-2},$
where $\{\textsf{X}_{2m-1}^{\ast},\,\textsf{Y}_{2m-1}^{\ast},\,\textsf{H}_{2m}^{\ast},\mid m\in\operatorname{\mathbb{Z}}\}$, such that
$\textsf{X}_{2m-1}^{\ast}=z^{m}\textsf{X}^{\ast},\quad\textsf{Y}_{2m-1}^{\ast}=z^{m-1}\textsf{Y}^{\ast},\quad\textsf{H}_{2m}^{\ast}=z^{m}\textsf{H}^{\ast},$
@@
$\textsf{H}^{\ast}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix},\quad\textsf{X}^{\ast}=\begin{pmatrix}0&0\\ 1&0\end{pmatrix},\quad\textsf{Y}^{\ast}=\begin{pmatrix}0&1\\ 0&0\end{pmatrix}.$
@@
Note, that $\mathfrak{g}_{-1}^{\ast}=\mathfrak{g}_{-1}$, and $\mathfrak{g}_{\ell}^{\ast}$ is dual to $\mathfrak{g}_{-\ell-2}$, if $\ell>-1$.